### v1

@@ 7.1   Basic setting
Taking $n=100,300,500,$ we choose two specifications to generate $y$ with $d_{\delta}=1$:
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$ (7.1)
@@
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},i=1,\ldots,n,$ (7.2)
[…]
Kelejian and Prucha (2007) consider the kernel function with a distance measure $d_{ij,m}^{*}$ and a threshold $d_{m}$. In estimation, we use the Epanechnikov kernel function, we choose the Euclidean distance $d_{ij}$ mentioned in $W_{1}$ to represent $d_{ij,m}^{*}$, and $d_{m}$ is considered as the quantile value of 10% of all distances. If $d_{ij,m}^{*}>d_{m}$, $\mathscr{K}\left(\frac{d_{ij,m}^{*}}{d_{m}}\right)=0$. We also consider $d_{\lambda}=2$ with $W_{1}$, $W_{2}$, and $\lambda=(0.5,0.3)^{\prime}$. The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[W_{1}X,W_{2}X,X,\Psi]$. We calculate the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%. We consider 1000 Monte Carlo replications for all experiments and $B=150$ bootstrap replications. We define $\mathscr{S}_{n}$ as the test statistic corresponding to $\mathbb{W}_{i}$ for $i=1,\ldots,4$ and $\mathscr{S}^{\ast}_{n}$ as the bootstrap version of $\mathscr{S}_{n}$.
The bootstrap version of our test, e.g., focusing on the setting described in Section 3, can be carried out by taking these steps:
  • Compute $\widehat{\epsilon}=y-L\hat{\xi}$ under null or alternative and obtain $\widetilde{\epsilon}=(I_{n}-\frac{1}{n}l_{n}l_{n}^{\prime})\widehat{\epsilon}$, where $l_{n}$ is the $n\times 1$ column vector of ones.
  • Sample randomly with replacement $n$ times from elements of $\widetilde{\epsilon}$ to obtain $\tilde{\epsilon}^{*}$.
  • Generate the bootstrap sample $\hat{y}=\left(I_{n}-\sum\limits_{j=1}^{d_{\lambda}}\hat{\lambda}_{j}W_{j}\right)^{+}\left([X,\Psi](\hat{\beta},\hat{\alpha})^{\prime}+\tilde{\epsilon}^{*}\right)$, where $(\cdot)^{+}$denotes the Moore-Penrose generalized inverse.
  • For the bootstrap, compute $\hat{\xi}^{\ast}$ and obtain the bootstrap residual $\hat{\epsilon}^{\ast}=\hat{y}-\hat{L}\hat{\xi}^{\ast}$ to construct $\mathscr{S}_{n}^{\ast}$, where $\hat{L}=[W_{1}\hat{y},\ldots,W_{d_{\lambda}}\hat{y},X,\Psi]$.
  • Repeat $B$ times to obtain the sequence of bootstrap test statistics $\left\{\mathscr{S}_{nj}^{\ast}\right\}_{j=1}^{B}$. Calculate bootstrap p-values $\mathscr{T}_{n}^{\ast}=\frac{1}{B}\sum\limits_{j=1}^{B}\mathbf{1}\left(\mathscr{S}_{n}<\mathscr{S}_{nj}^{\ast}\right)$ and reject the null hypothesis if $\mathscr{T}_{n}^{\ast}$ is smaller than the prescribed nominal level of significance.
As before, we denote by $\mathscr{T}_{n}$ the p-value based on the asymptotic test. From Tables 14-15, we find that the bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases. Specifically, we observe that the pairs $(n,h)=(500,8)$ deliver acceptable sizes under three disturbances. Moreover, we observe that the asymptotic test is powerful against local alternatives as $n$ increases from Tables S1-S2 in the supplement.
We also consider the case of $d_{\lambda}=2$ in the supplement. The size performances [^5] Power performances are close to or equal to 1 in most cases so we omit the power tables but they are available on request. are reported in Tables S3-S4 in the supplement and the finite sample properties are similar to the case of $d_{\lambda}=1$.
@@ 7.2   Nonparametric spatial matrix
For the nonparametric spatial matrix, we generate $w_{1ij}=\Phi\left(-b_{ij}\right)\mathbf{1}\left(c_{ij}<0.1\right)$ if $i\neq j$, and $w_{1ii}=0$, where $\Phi(\cdot)$ is the standard normal cdf, $b_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-3,3]$, and $c_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. From this construction, we ensure that $W_{1}$ is sparse with no more than $10\%$ elements being nonzero. Then, $W=W_{1}/1.2\bar{\alpha}(W_{1})$, ensuring the existence of $(I-W)^{-1}$. For $\mathbb{W}_{3}$, we approximate elements in $W$ by $\widehat{w}_{ij}=\sum_{l=0}^{r}\tau_{l}[d_{ij,m}^{*}]^{l}\mathbf{1}(d_{ij,m}^{*}<d_{m})$ if $i\neq j$ with $r=2,4$, where $d_{ij,m}^{*}$ and $d_{m}$ are discussed in Section 7.1, and the IV matrix for $\mathbb{W}_{3}$ is $K_{2}=[X,\Psi]$. We define this DGP setting as (D3).
Table 16 reports the simulation results at the 1%, 5% and 10% asymptotic level with $r=2,4$ and $h=2,4,8$. We find that there is a clear pattern of rejection rates approaching the theoretical level as sample size increases, and sizes are acceptable for $n=500$ and $h=8$. When $r$ increases, the results are similar. From Table S5, the power performances are close to or equal to 1 as $n$ increases. We also consider $W=W_{1}+W_{2}$ which corresponds to $d_{\lambda}=2$. Specifically, we set $w_{ij2}=d_{m}\exp(-d_{ij,m}^{*}/d_{m})$ if $d_{ij,m}^{*}\leq d_{m}$ and $w_{ij2}=0$ if $d_{ij,m}^{*}>d_{m}$. From Table S6 in the supplement, a similar pattern for size is observed.
@@ 7.3   Varying spatial coefficient
We now consider varying $\lambda$ such that $\lambda=\lambda(z_{i})$. When $d_{\lambda}=1$, we set $\lambda(z_{i})=\sin(0.75\pi\times z_{i})$, when $d_{\lambda}=2$, we set $\lambda_{1}(z_{i})=\sin(0.38\pi\times z_{i})$ and $\lambda_{2}(z_{i})=\cos(0.37\pi\times z_{i})$. Others are same as those in Section 7.1. For estimation, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$. Then, $h_{1il}=y_{i}\sum_{j=1}^{n}w_{1ij}\phi_{jl}(z)$ and $h_{2il}=y_{i}\sum_{j=1}^{n}w_{2ij}\phi_{jl}(z)$. The IV matrix is $K_{3}=[H_{1}X,X,\Psi]$ for $d_{\lambda}=1$ and $K_{3}=[H_{1}X,H_{2}X,X,\Psi]$ for $d_{\lambda}=2$. We define this DGP setting as (D4).
Tables 17 and 18 report the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%. Sizes are acceptable as $n$ and $h$ increase, especially, the trigonometric approximation performs slightly better than polynomial approximations for Chi-square distribution when testing $\lambda=0$. Bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases. When $d_{\lambda}=2$, the results are similar, as shown in Tables S7-S8 in the supplement.
From Tables S9 and S10 in the supplement, power improves as $h$ increases for all designs, and the test for $\lambda$ and $\delta$ have similar results. As for $d_{\lambda}=2$, we find that our proposed test statistics still perform well as both $n$ and $h$ increase.
7.4   Additional simulation results in Online Supplement
In a range of experiments reported in the online supplement, we study how large $n$ with $h$ increasing to see how size changes. So we set $n=1000$ with $h=12,16$, $d_{\lambda}=1,2$ and $d_{\delta}=1$. From Table S11 of the supplement, we find that as $h$ increases, the performance of size improves.
We also consider tests of equality of varying coefficients and CRS hypothesis, where $d_{\delta}=2$ such that
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+\epsilon_{i},$ (7.3)
and
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},$ (7.4)
for $i=1,\ldots,n$. Our null of interest when testing equality is $H_{0}^{true}:\delta_{1}(z)=\delta_{2}(z),z\in\mathcal{Z},$ while for CRS it is $H_{0}^{true}:\delta_{1}(z)+\delta_{2}(z)=1,z\in\mathcal{Z}.$ In the first case, we reparameterize with $\theta(z)=\delta_{1}(z)-\delta_{2}(z)$, and the model is $y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i1}+p_{i2})\delta_{2}(z_{i})+\epsilon_{i}.$ In the second case, we reparameterize with $\theta(z)=\delta_{1}(z)+\delta_{2}(z)-1$ and the model is $y_{i}-p_{i1}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i2}-p_{i1})\delta_{2}(z_{i})+\epsilon_{i}.$
The results are included in the supplement. Taking CRS as an example, from Tables S12-S13, bootstrap tests are closer to the nominal levels than the asymptotic tests although the sizes of both types of tests improve generally as the sample size increases. For the nonparametric spatial matrix, from Table S14, there is a clear pattern of rejection rates approaching the theoretical level as sample size increases. The results of testing equality are similar, as shown in Tables S15-S17.
@@ Appendix A General Central Limit Theorem
where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7.
(b) The entries $s_{ij}$ of the symmetric matrix $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$.
(c) $\mathscr{S}$ satisfies 1.
@@
[…]

### v2

@@ 7.1 Basic setting
Taking $n=200,500,900,$ we choose two specifications to generate $y$ with $d_{\delta}=1$:
$y_{i}=\sum_{k=1}^{d_{\lambda}}\lambda_{k}\sum\limits_{j=1}^{n}w_{kij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$
@@
$y_{i}=\sum_{k=1}^{d_{\lambda}}\lambda_{k}\sum\limits_{j=1}^{n}w_{kij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+v_{i},\ v_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}+\epsilon_{i},i=1,\ldots,n,$
where $\epsilon_{i}$ are $i.i.d$ with three zero mean and unit variance distributions: (V1) $N(0,1)$, (V2) $\sqrt{\frac{5}{4}}t(10)$ and (V3) $\frac{1}{4}(\chi_{8}^{2}-8)$. We generate $z_{i}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$, $p_{i}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-2,2]$, $x_{i1}=1$, $x_{i2}\stackrel{{\scriptstyle i.i.d}}{{\sim}}N(1,2)$ and set $\beta=(-1,1)^{\prime}$. For $\lambda$, we fix the total weight at 0.9 and then define a decreasing vector $\lambda^{*}=(d_{\lambda},d_{\lambda}-1,...,1)^{\prime}$. We then set $\lambda=0.9\left(\sum_{k=1}^{d_{\lambda}}\lambda_{k}^{*}\right)^{-1}\lambda^{*}$.
For the spatial weights matrices, we generate $W_{k}$ as circulants. Specifically, let $W_{k}^{*}$ be the symmetric circulant matrix whose first-row entries are $w_{k1j}^{*}=0$ if $j=1$ or $j=k+2,\ldots,n-k$ and $w_{k1j}^{*}=1$ if $j=2,\ldots,k+1$ or $j=n-k+1,\ldots,n$. Thus, the weight matrix $W_{k}^{*}$ encapsulates a binary neighbourhood criterion for $k$ neighbours on either `side' of a unit. Now define the normalized matrix $W_{k}=\{\overline{\alpha}(W_{k}^{*})\}^{-1}W_{k}^{*},$ recalling that $\overline{\alpha}(W_{k}^{*})=2k$ for a circulant matrix. Furthermore, $b_{ij}$ is the typical element of $(I_{n}-\sum_{k=1}^{d_{\lambda}}\lambda_{k}W_{k})^{-1}$.
To construct the SHAC statistics $\mathbb{W}_{2},\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we proceed as follows. Generate $M=d_{\lambda}$ distinct distance measures by setting $\ell=[n^{\eta}]+1$ with $\eta=3/7$ and taking $q=8$, so that each unit has at least $\ell$ neighbors, where $\eta$ is defined in Assumption 9 (a). [^4] From the symmetric circulant matrix $W_{d_{\lambda}}$, we build a graph and compute its pairwise shortest‐path distances in MATLAB, then apply cmdscale to generate $n\times 2$ coordinate matrix. We next compute the full $n\times n$ Euclidean distance matrix $D$ with the typical elements $d_{ij}$. For each m=1,,Mm=1,...,M, we add the actual distance matrix, $D_{m}^{*}$, used in practice with unobserved symmetric measurement noise. Its typical entries are $d_{ij,m}^{*}=d_{ij}+\nu_{ij,m},$ where $\nu_{ij,m}=1/2(\mu_{ij,m}+\mu_{ji,m})$ with $\mu_{ij,m}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. Define for each unit $i$ the $\ell$-th nearest‐neighbor distance by $d_{(i,\ell),m}=$ the $\ell$-th smallest entry of row $i$ of $D^{*}_{m}$, excluding the diagonal. We then set $d_{m}=\max_{1\leq i\leq n}d_{(i,\ell),m},$ so that every unit has at least $\ell$ neighbors within distance $d_{m}$. Collecting these over $m=1,\dots,M$ yields the $M\times 1$ vector $d_{m}$. We set $\delta=0$ for the null hypothesis and $\delta(\mathrm{x})=1-\mathrm{x}^{2}$ for any $\mathrm{x}\in R$ for the alternatives. We choose polynomials $\psi_{j}(\mathrm{x})=\mathrm{x}^{j}$ for $j=1,\cdots,h$ and the design of trigonometric functions is in the supplement. The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[X,W_{1}X,...,W_{d_{\lambda}}X,\Psi]$. Furthermore, asy-p represents the standard-normal p-values, whereas chi-p represents the chi-square calibration $(\chi^{2}_{d_{\alpha}}-d_{\alpha})/\sqrt{2d_{\alpha}}$, which is to improve testing performance in small samples. This may also help in empirical applications, although in both applications in the previous section critical values from either distribution yield the same results. We discuss size performance in this section and present power performance in the supplement, since under global alternatives the power of the asymptotic tests tends to one as $n\to\infty$.
Table 4 reports the empirical sizes of $\mathbb{W}_{1}$, and Table 5 reports those of the SHAC-corrected statistic $\mathbb{W}_{2}$, at nominal 1%, 5% and 10% levels. When $n=200$, $\mathbb{W}_{2}$ exhibits substantially larger size distortion than $\mathbb{W}_{1}$. The reason is: Theorem 2.1 requires $n^{-1}\,d_{\xi}\to 0,$ whereas Theorem 3.1 requires $n^{\eta-1}\,d_{\xi}\to 0$ with $\eta=3/7$. In our design $d_{\xi}=d_{\lambda}+d_{\beta}+d_{h}$ with $d_{\lambda}\in\{2,4\}$, $d_{\beta}=2$, $d_{h}\in\{2,4,8\}$, so $d_{\xi}/n$ is already small at $n=200$, but $n^{\eta-1}d_{\xi}^{2}$ is relatively large. As $n$ increases to 500 and 900, $n^{\eta-1}d_{\xi}^{2}$ shrinks and the size of $\mathbb{W}_{2}$ converges rapidly to its nominal levels. Table 4-5 also show a clear pattern across the polynomial order with $h\in\{2,4,8\}$. For both $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$, the smallest sieve yields the most accurate sizes, the largest sieve the greatest distortion, particularly for $\mathbb{W}_{2}$ at $n=200$. This accords with the fact that larger $h$ increases $d_{\alpha}$ and hence $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$, which in turn slows the convergence rates $n^{-1}d_{\xi}\to 0$ for $\mathbb{W}_{1}$ and $n^{\eta-1}d_{\xi}^{2}\to 0$ for $\mathbb{W}_{2}\,$. A similar scenario obtains when increasing $d_{\lambda}$, since it also raises $d_{\xi}$ and worsens finite-sample size accuracy.
@@ 7.2 Nonparametric spatial matrix
For the nonparametric spatial matrix, we generate $g^{*}_{ij}=\Phi\left(-b_{ij}\right)\mathbf{1}\left(c_{ij}<0.1\right)$ if $i\neq j$, and $w_{1ii}=0$, where $\Phi(\cdot)$ is the standard normal cdf, $b_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-3,3]$, and $c_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. From this construction, we ensure that $G^{*}$ is sparse with no more than $10\%$ elements being nonzero. Then, define $G=G^{*}/1.2\bar{\alpha}(G^{*})$, ensuring the existence of $(I-G)^{-1}$. For $\mathbb{W}_{3}$, we approximate elements in $G$ by $\widehat{g}_{ij}=\sum_{l=0}^{d_{\tau}}\tau_{l}e_{l}(d_{ij})=\sum_{l=0}^{d_{\tau}}\tau_{l}[d_{ij,m}^{*}]^{l}\mathbf{1}(d_{ij,m}^{*}<d_{m})$ if $i\neq j$ with $d_{\tau}=2,4$, where $d_{ij,m}^{*}$ and $d_{m}$ are discussed in Section 7.1, and the IV matrix for $\mathbb{W}_{3}$ is $K_{2}=[X,\mathcal{E}_{1}X,...,\mathcal{E}_{d_{\tau}}X,\Psi]$, where $e_{l}(d_{ij})$ are the typical elements of $\mathcal{E}_{l}$.
Table 6 reports the empirical sizes of $\mathbb{W}_{3}$ at nominal 1%, 5% and 10% levels. When $n=200$, $\mathbb{W}_{3}$ with $d_{\tau}=2$ yields rejection rates very close to the nominal levels, whereas the larger basis $d_{\tau}=8$ produces mild over-rejection. As $n$ increases to 500 and 900, both choices of $d_{\tau}$ rapidly converge to their nominal sizes. This scenario is similar to that of $\mathbb{W}_{2}$ and originates from the rate condition $n^{\eta-1}d_{\tau}^{2}\to 0$ in Theorem 4.1. In other words, it also requires an accurate approximation of the SHAC estimator.
@@ 7.3 Varying spatial coefficient
We now consider varying $\lambda$ such that $\lambda=\lambda(z_{i})$ and $d_{\lambda}=1,2$. When $d_{\lambda}=1$, we set $\lambda(z_{i})=0.9\sin(\pi z_{i})$, when $d_{\lambda}=2$, we set $\lambda_{1}(z_{i})=0.6\sin(\pi z_{i})$ and $\lambda_{2}(z_{i})=0.3\sin(\pi z_{i})$. Others are the same as those in Section 7.1. For estimation, we use $\phi_{il}(z)=\frac{1}{h}[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\frac{1}{h}\sin(\frac{z_{i}}{2l})$ as the trigonometric function for l=1,,hl=1,...,h. The IV matrix $K_{3}=K_{1}$ with $d_{\lambda}=1,2$.
Table 7 reports the rejection probabilities of $\mathbb{W}_{4}$ at nominal 1%, 5% and 10% levels when testing $\delta(z)$. At $n=200$, the smallest sieve $(h=2,\,d_{\lambda}=1)$ achieves sizes close to nominal, whereas larger sieves, especially $h=8,\,d_{\lambda}=2$, over‐reject noticeably. As $n$ increases to 900, size distortion vanishes rapidly. Table S1 shows the sizes of $\mathbb{W}_{4}$ when testing $\lambda(z)$. In particular, at $n=200$ with $d_{\lambda}=2$ and $h=8$, the over‐rejection is especially severe. By $n=900$, size accuracy is largely restored. This pattern reflects the rate condition in Theorem 5.1, $n^{\eta-1}d_{\gamma}^{2}\to 0$, since both $h$ and $d_{\lambda}$ enter the total sieve dimension $d_{\gamma}$, slowing the approximation when they are large. Across all designs, the chi-p delivers empirical sizes closer to nominal levels than asy-p in small samples, while the difference dissipates as $n$ grows. Results for trigonometric functions are similar, and left for the online appendix.
8.  Conclusion
We provide a machinery for conducting tests on nonparametrically varying coefficients using ideas from parametric testing, and with similar ease of implementation. We allow for spatial dependence and provide for testing nonparametrically varying spatial dependence coefficients. It seems reasonable to conjecture that the methodology is applicable in other settings where one may be interested in testing linear restrictions on nonparametric functions. These provide directions for future research.
@@ Appendix A General Central Limit Theorem
where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7 and $\mathscr{S}$ is a symmetric matrix. (b) The entries $s_{ij}$ of $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$. (c) $\mathscr{S}$ satisfies 1.
@@
${\left(v^{\prime}\mathscr{A}v-J\right)}/{\sqrt{2J}}\overset{d}{\longrightarrow}N(0,1).$ (A.2)